OpenStudy (anonymous):

2x - y - 4 = 0 3x + y - 9 = 0 What is the solution set of the given system?

OpenStudy (anonymous):

2x-y=4 3x+y=9 The ys cancel out and you add your xs and constants. 5x=13 x=13/5 Take the x value and plug it into one of the equations to find your y.

OpenStudy (anonymous):

{[13/5 6/5]}

OpenStudy (anonymous):

?

OpenStudy (anonymous):

good!